# Finite-Agent Stochastic Differential Games on Large Graphs: I. The Linear-Quadratic Case

### v1

@@ 4 Semi-Explicit Equilibrium under Vertex-Transitive Graph
$\text{Aut}(G):=\{\text{bijection }\phi:V\to V|(u,v)\in E\text{ if and only if }(\phi(u),\phi(v))\in E,\ \forall u,v\in V\}.$ (4.1)
@@
where the inequality between symmetric matrices is defined in terms of the semi-positive-definite sense. That is, $A\geq B$ if and only if $A-B$ is semi-positive-definite, for $A,B\in\mathbb{S}^{N\times N}$.
@@ 4.1 Main results
$R^{\prime}(t)=\frac{1}{c}\emph{\text{Tr}}\left[Q^{\prime}(R(t))e^{-t(a+q)L}\right]e^{-t(a+q)L},\quad R(0)=0,$ (4.3)
@@
$F^{i}_{t}=\frac{1}{\frac{\emph{\text{Tr}}(P_{t})-(a+q)\emph{\text{Tr}}(L)}{N}}[P_{t}-(a+q)L]e_{i}e_{i}^{\operatorname{T}}[P_{t}-(a+q)L],$ (4.5)
@@
where $P_{t}$ is defined in equation (4.20) and has the representation (4.6) given by Theorem 4.7. Given the initial condition $\hat{X}_{0}=X_{0}$, this Ornstein-Uhlenbeck dynamics has a closed-form solution:
@@
where we have used the fact that $P_{s}$ and $P_{t}$ commute for any $t,s\in[0,T]$. Therefore, $\hat{X}_{t}$ is Gaussian distributed with mean $e^{-\int_{0}^{t}P_{s}\,\mathrm{d}s}X_{0}$ and variance $\sigma^{2}\int_{0}^{t}e^{-2\int_{s}^{t}P_{u}\,\mathrm{d}u}\,\mathrm{d}s$. Using equations (4.35) and (4.53) in Section 4.2 below, which can be verified to be correct, the mean and variance of $\hat{X}_{t}$ can be represented in terms of $R$, the solution to equation (4.3).
We note the following, (i) Theorem 4.7 applies specifically to the case where $q^{2}=\varepsilon$. As will discussed in Remark 4.8, this condition results in an uncoupled nature of the underlying linear system, greatly simplifying the calculations. (ii) A numerical verification of Theorem 4.7 will be presented in Section 5.1. (iii) Theorem 4.7 can be generalized to provide a semi-explicit Markovian NE for games where the mean reversion level is the average of states within a more broadly defined neighborhood of a player, such as the $l$-neighborhood (see Definition 4.12 and Remark 4.19). This extends the concept of a graph neighborhood as introduced in Definition 2.1. The generalized result is briefly outlined below.
@@
Remark 4.5 (Generalizations of Theorem 4.7).
To set the the mean reversion level as the average within the $l$-neighborhood, we consider the following generalized state dynamics on a vertex-transitive graph $G$:
@@
$f^{i}(t,x,\alpha)=\frac{1}{2}(\alpha)^{2}-q\alpha\left(\sum_{j:v_{j}\in N_{G}^{l}(v_{i})}\frac{n(v_{j},v_{i};l)}{|N_{G}^{l}(v_{i})|}x^{j}-x^{i}\right)+\frac{\varepsilon}{2}\left(\sum_{j:v_{j}\in N_{G}^{l}(v_{i})}\frac{n(v_{j},v_{i};l)}{|N_{G}^{l}(v_{i})|}x^{j}-x^{i}\right)^{2},$ (4.14)
$g^{i}(x)=\frac{c}{2}\left(\sum_{j:v_{j}\in N_{G}^{l}(v_{i})}\frac{n(v_{j},v_{i};l)}{|N_{G}^{l}(v_{i})|}x^{j}-x^{i}\right)^{2}.$ (4.15)
@@
where each vertex $v_{j}\in\bigcup_{k=1}^{l}N^{k}_{G}(v_{i})$ is assigned a weight $\sum_{k=1}^{l}w^{k}n(v_{j},v_{i};k)/\sum_{k=1}^{l}w^{k}|N_{G}^{k}(v_{i})|$, proportional to a discounted sum of the number of paths of length $k\in\left\{1,\ldots,l\right\}$ between $v_{i}$ and $v_{j}$. Each path of length $k$ is discounted by a factor of $w^{k}$, where $w$ denotes the discount rate.
The expression in equation (4.18) can be written in matrix form:
$-e_{i}^{\operatorname{T}}S(l,L)X_{t},\quad\text{ where}\quad S(l,L)=I-\frac{1-w\delta}{1-(w\delta)^{l}}(I-L)[I-(w\delta)^{l}(I-L)^{l}][I-w\delta(I-L)]^{-1}.$ (4.19)
@@
4.2 Heuristic derivations
@@
To construct the solution in equation (4.5), we draw inspiration from a previous work [LS22] and employ two key techniques: a fixed point scheme that aids in representing the solution to the Riccati system and the regular representation of the graph automorphism group that leverages symmetry. We will discuss them below one by one.
During the heuristic derivations below, several assumptions must be made. For a smoother presentation flow, these assumptions (A.1)–(A.4) are summarized later in Remark 4.11, and are referred to during the heuristic derivation. However, we emphasize that verifying these assumptions is not necessary due to the existence of the verification procedure presented in Section 4.4.
Let $P:[0,T]\to\mathbb{R}^{N\times N}$ be a matrix-valued function defined in terms of $F^{1}_{t},\cdots,F^{N}_{t}$:
$P_{t}:=\sum_{k=1}^{N}\left[(a+q)L+F^{k}_{t}\right]e_{k}e_{k}^{\operatorname{T}}=(a+q)L+\sum_{k=1}^{N}F^{k}_{t}e_{k}e_{k}^{\operatorname{T}}.$ (4.20)
$\dot{F}^{i}_{t}-P_{t}F^{i}_{t}-F^{i}_{t}P_{t}+(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L+F^{i}_{t}e_{i}e_{i}^{\operatorname{T}}F^{i}_{t}=0,\quad F^{i}_{T}=cLe_{i}e_{i}^{\operatorname{T}}L.$ (4.21)
Treating $P$ as given, the solution to equation (4.21) can be represented in terms of $P_{t}$, denoted as $F^{1}_{t}(P),\cdots,F^{N}_{t}(P)$. Substituting $F^{1}_{t}(P),\cdots,F^{N}_{t}(P)$ into equation (4.20) provides an equation solely for $P$:
$P_{t}=(a+q)L+\sum_{k=1}^{N}F^{k}_{t}(P)e_{k}e_{k}^{\operatorname{T}}.$ (4.22)
Definition 4.6.
$R_{\varphi}e_{i}=e_{\varphi(i)}.$ (4.23)
Immediately from Definition 4.23, one has $R_{\varphi}R_{\psi}=R_{\varphi\circ\psi}$, $R^{-1}_{\varphi}=R_{\varphi^{-1}}=R_{\varphi}^{\operatorname{T}}$, $\forall\varphi,\psi\in\text{Aut}(G)$. The following lemma outlines key properties for leveraging the symmetry of vertex-transitive graphs.
Lemma 4.7 ([LS22, Lemma 4.1]).
  1. (i)
    $L$ commutes with $R_{\varphi}$ for every $\varphi\in\emph{\text{Aut}}(G)$.
  2. (ii)
    If $Y\in\mathbb{R}^{N\times N}$ commutes with $R_{\varphi}$ for every $\varphi\in\emph{\text{Aut}}(G)$, then $Y_{ii}=\frac{1}{N}\emph{\text{Tr}}(Y)$, $\forall i\in[N]$.
  3. (iii)
    If $Y^{1},\cdots,Y^{N}\in\mathbb{R}^{N\times N}$ are such that $R_{\varphi}Y^{i}=Y^{\varphi(i)}R_{\varphi}$ holds for every $\varphi\in\emph{\text{Aut}}(G)$ and $i\in[N]$, then $Y_{ii}^{i}=Y_{jj}^{j}$, $\forall i,j\in[N]$.
4.2.1 Solving $F^{i}$ in terms of $P$
Our first step is to solve the Riccati equation (4.21) for $F_{t}^{i}$, treating $P$ as given, and represent the solution in terms of $P$. To this end, we consider the linear system for $(Y_{t}^{i},\Lambda_{t}^{i})$:
$\begin{bmatrix}\dot{Y}^{i}_{t}\\ \dot{\Lambda}^{i}_{t}\end{bmatrix}=\begin{bmatrix}P_{t}&-(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L\\ e_{i}e_{i}^{\operatorname{T}}&-P_{t}\end{bmatrix}\begin{bmatrix}Y^{i}_{t}\\ \Lambda^{i}_{t}\end{bmatrix},$ (4.24)
with terminal conditions $Y^{i}_{T}=cLe_{i}e_{i}^{\operatorname{T}}L$, $\Lambda^{i}_{T}=I$. If the solution to equation (4.24) is found and $\Lambda^{i}_{t}$ is invertible for any $t\in[0,T]$, then $F^{i}_{t}=Y^{i}_{t}(\Lambda^{i}_{t})^{-1}$ is the solution to equation (4.21) [BJP00]. Under the condition $q^{2}=\varepsilon$, the equation for $Y_{t}^{i}$ becomes decoupled, resulting in $Y^{i}_{t}=\tilde{P}_{t}Y^{i}_{T}$, where
[…]
By substituting $Y^{i}_{t}=\tilde{P}_{t}Y^{i}_{T}$ into equation (4.24), we obtain an ODE for $\Lambda_{t}^{i}$: $\dot{\Lambda}^{i}_{t}=e_{i}e_{i}^{\operatorname{T}}\tilde{P}_{t}Y^{i}_{T}-P_{t}\Lambda^{i}_{t}$. The solution to this equation: $\Lambda^{i}_{t}=\tilde{P}^{-1}_{t}\left(\Lambda^{i}_{T}-\int_{t}^{T}\tilde{P}_{s}e_{i}e_{i}^{\operatorname{T}}\tilde{P}_{s}Y^{i}_{T}\,\mathrm{d}s\right)$. Therefore, incorporating the terminal conditions, the underlying linear system yields the following solution:
$\begin{cases}Y_{t}^{i}=c\tilde{P}_{t}Le_{i}e_{i}^{\operatorname{T}}L\\ \Lambda_{t}^{i}=\tilde{P}^{-1}_{t}\left(I-c\int_{t}^{T}\tilde{P}_{s}e_{i}e_{i}^{\operatorname{T}}\tilde{P}_{s}Le_{i}e_{i}^{\operatorname{T}}L\,\mathrm{d}s\right)\end{cases}.$ (4.26)
Assuming (A.2) and applying Lemma 4.7(ii) gives,
$e_{i}^{\operatorname{T}}\tilde{P}_{s}Le_{i}=\frac{1}{N}\text{Tr}(\tilde{P}_{s}L).$ (4.27)
This simplifies $\Lambda_{t}^{i}$ to:
$\Lambda_{t}^{i}=\tilde{P}^{-1}_{t}\left(I-c\int_{t}^{T}\frac{\text{Tr}(\tilde{P}_{s}L)}{N}\tilde{P}_{s}e_{i}e_{i}^{\operatorname{T}}L\,\mathrm{d}s\right).$ (4.28)
Under assumption (A.3), the solution to equation (4.21) is provided as
$F^{i}_{t}=c\tilde{P}_{t}Le_{i}e_{i}^{\operatorname{T}}L\left(I-c\int_{t}^{T}\frac{\text{Tr}(\tilde{P}_{s}L)}{N}\tilde{P}_{s}e_{i}e_{i}^{\operatorname{T}}L\,\mathrm{d}s\right)^{-1}\tilde{P}_{t}.$ (4.29)
Up until now, we’ve derived a solution $F^{i}_{t}$ expressed in terms of $P$.
Remark 4.8.
The condition $q^{2}=\varepsilon$ ensures that the underlying linear system for $(Y^{i}_{t},\Lambda^{i}_{t})$ is decoupled. Without this condition, we need to solve the full system (4.24), for which the existence of a closed-form solution remains unclear since:
$\begin{bmatrix}P_{t}&-(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L\\ e_{i}e_{i}^{\operatorname{T}}&-P_{t}\end{bmatrix}\text{ does not commute with }\begin{bmatrix}P_{s}&-(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L\\ e_{i}e_{i}^{\operatorname{T}}&-P_{s}\end{bmatrix}\text{ when }t\neq s.$
4.2.2 Expressing $P$ in terms of $\eta$
To further simplify equation (4.29), we define
$\Sigma^{i}_{t}:=\int_{t}^{T}\frac{\text{Tr}(\tilde{P}_{s}L)}{N}\tilde{P}_{s}e_{i}e_{i}^{\operatorname{T}}L\,\mathrm{d}s,$ (4.30)
and
$\eta^{i}_{t}:=e_{i}^{\operatorname{T}}L\left(I-c\Sigma^{i}_{t}\right)^{-1}\tilde{P}_{t}e_{i}.$ (4.31)
Thus $F_{t}^{i}$ admits the representation $F^{i}_{t}=c\tilde{P}_{t}Le_{i}e_{i}^{\operatorname{T}}L\left(I-c\Sigma^{i}_{t}\right)^{-1}\tilde{P}_{t}$ and satisfies $F^{i}_{t}e_{i}=c\eta^{i}_{t}\tilde{P}_{t}Le_{i}$. Under assumption (A.2), using Lemma 4.7(i) and equation (4.30) leads to $R_{\varphi}\Sigma^{i}_{t}=\Sigma^{\varphi(i)}_{t}R_{\varphi}$, $\forall\varphi\in\text{Aut}(G)$. With the additional assumption (A.4), we find:
$R_{\varphi}L\left(I-c\Sigma^{i}_{t}\right)^{-1}=L\left(I-c\Sigma^{\varphi(i)}_{t}\right)^{-1}R_{\varphi},\ \forall\varphi\in\text{Aut}(G).$ (4.32)
By Lemma 4.7(iii), $\eta^{i}_{t}$ becomes independent of the player index $i$. From this point, we will use the notation $\eta_{t}$ instead of $\eta_{t}^{i}$, as this quantity is shared by all players.
We now proceed to derive an expression for $P_{t}$ in terms of $\eta_{t}$. Substituting the relation $F^{i}_{t}e_{i}=c\eta_{t}\tilde{P}_{t}Le_{i}$ into equation (4.20) yields
$P_{t}=(a+q)L+c\eta_{t}\tilde{P}_{t}L,$ (4.33)
where $\tilde{P}$ is defined as in equation (4.25). It remains to solve the above equation for $P_{t}$.
Under assumption (A.1), any two symmetric matrices in the collection $\{L\}\cup\{P_{t},\tilde{P}_{t}:t\in[0,T]\}$ commute, which implies that this collection of matrices can be diagonalized simultaneously. Therefore, one can work with the spectra of the matrices rather than directly with the matrices themselves, providing greater convenience for calculations. Let $\rho^{1}_{t},\ldots,\rho^{N}_{t}$ be the eigenvalues of $P_{t}$ and $\lambda_{1},\ldots,\lambda_{N}$ be the eigenvalues of $L$, equation (4.33) becomes
$\rho^{k}_{t}=(a+q)\lambda_{k}+c\eta_{t}\lambda_{k}e^{-\int_{t}^{T}\rho^{k}_{s}\,\mathrm{d}s},\ \forall k\in[N].$ (4.34)
$e^{\int_{t}^{T}\rho^{k}_{s}-(a+q)\lambda_{k}\,\mathrm{d}s}=1+c\lambda_{k}\int_{t}^{T}\eta_{s}e^{-(T-s)(a+q)\lambda_{k}}\,\mathrm{d}s.$ (4.35)
$\rho^{k}_{t}=(a+q)\lambda_{k}+\frac{c\lambda_{k}\eta_{t}e^{-(T-t)(a+q)\lambda_{k}}}{1+c\lambda_{k}\int_{t}^{T}\eta_{s}e^{-(T-s)(a+q)\lambda_{k}}\,\mathrm{d}s}.$ (4.36)
$P_{t}=(a+q)L+c\eta_{t}Le^{-(T-t)(a+q)L}\left(I+\int_{t}^{T}c\eta_{s}e^{-(T-s)(a+q)L}\,\mathrm{d}s\;L\right)^{-1}.$ (4.37)
The above expression for $P_{t}$ in terms of $\eta_{t}$ is not entirely satisfactory because $\eta$, as defined in (4.31), depends on $\tilde{P}_{t}$, and therefore on $P_{t}$. Consequently, equation (4.37) does not provide an explicit construction for $P_{t}$. The next step (in Section 4.2.3) is to find a way to characterize and construct $\eta$ based solely on model parameters. Once this is achieved, equation (4.37) can be used to construct $P_{t}$, as detailed in Section 4.2.4.
4.2.3 The governing equation for $\eta$
According to the definition of $\eta$ in (4.31), the primary challenge in characterizing $\eta_{t}$ arises from the inverse term $\left(I-c\Sigma^{i}_{t}\right)^{-1}$. Combining equations (4.27) and (4.30) gives
$(\Sigma^{i}_{t})^{2}=\left[\int_{t}^{T}\left(\frac{\text{Tr}(\tilde{P}_{s}L)}{N}\right)^{2}\,\mathrm{d}s\right]\Sigma^{i}_{t}.$ (4.38)
$\left(I-c\Sigma^{i}_{t}\right)^{-1}=I+c\left[1-c\int_{t}^{T}\left(\frac{\text{Tr}(\tilde{P}_{s}L)}{N}\right)^{2}\,\mathrm{d}s\right]^{-1}\Sigma^{i}_{t}.$ (4.39)
Plugging it back into equation (4.31) and combining it with equation (4.27) yields
$\eta_{t}=\frac{g_{t}}{1-c\int_{t}^{T}g^{2}_{s}\,\mathrm{d}s},$ (4.40)
where $g_{t}$ is a scalar function defined as:
$g_{t}:=\frac{\text{Tr}(\tilde{P}_{t}L)}{N}.$ (4.41)
A property of $g_{t}$ will be repeatedly referred to later, so we derive it here. Direct calculations based on equation (4.33) gives:
$\text{Tr}(\tilde{P}_{t}L)=\frac{1}{c\eta_{t}}\left[\text{Tr}(P_{t})-(a+q)\text{Tr}(L)\right].$ (4.42)
Given $\text{Tr}(P_{t})=\sum_{k=1}^{N}\rho^{k}_{t}$, combining it with equation (4.36) yields:
$Ng_{t}=\text{Tr}(\tilde{P}_{t}L)=\sum_{k=1}^{N}\frac{\lambda_{k}e^{-(T-t)(a+q)\lambda_{k}}}{1+c\lambda_{k}\int_{t}^{T}\eta_{s}e^{-(T-s)(a+q)\lambda_{k}}\,\mathrm{d}s}.$ (4.43)
Back to equation (4.40), by multiplying both sides by $cg_{t}$, integrating in $t$, and taking the exponential, we obtain:
$e^{c\int_{t}^{T}g_{s}\eta_{s}\,\mathrm{d}s}=e^{\int_{t}^{T}\frac{cg^{2}_{s}}{1-c\int_{s}^{T}g^{2}_{u}\,\mathrm{d}u}\,\mathrm{d}s}=\frac{1}{1-c\int_{t}^{T}g^{2}_{s}\,\mathrm{d}s}.$ (4.44)
Combining equations (4.41) and (4.43), one has
$c\int_{t}^{T}g_{s}\eta_{s}\,\mathrm{d}s=\frac{1}{N}\sum_{k=1}^{N}\int_{t}^{T}\frac{c\lambda_{k}\eta_{s}e^{-(T-s)(a+q)\lambda_{k}}}{1+c\lambda_{k}\int_{s}^{T}\eta_{u}e^{-(T-u)(a+q)\lambda_{k}}\,\mathrm{d}u}\,\mathrm{d}s.$ (4.45)
Combining equations (4.44)–(4.45) and using similar arguments as done for equation (4.35) implies
$\frac{1}{1-c\int_{t}^{T}g^{2}_{s}\,\mathrm{d}s}=\left[\prod_{k=1}^{N}\left(1+c\lambda_{k}\int_{t}^{T}\eta_{s}e^{-(T-s)(a+q)\lambda_{k}}\,\mathrm{d}s\right)\right]^{\frac{1}{N}}.$ (4.46)
Plugging equations (4.41), (4.43) and (4.46) into equation (4.40) yields
$\eta_{t}=\frac{1}{N}\left(\sum_{k=1}^{N}\frac{\lambda_{k}e^{-(T-t)(a+q)\lambda_{k}}}{1+c\lambda_{k}\int_{t}^{T}\eta_{s}e^{-(T-s)(a+q)\lambda_{k}}\,\mathrm{d}s}\right)\left[\prod_{k=1}^{N}\left(1+c\lambda_{k}\int_{t}^{T}\eta_{s}e^{-(T-s)(a+q)\lambda_{k}}\,\mathrm{d}s\right)\right]^{\frac{1}{N}}.$ (4.47)
Finally, rewriting equation (4.47) in the matrix form we obtain:
$\eta_{t}=\frac{1}{N}\text{Tr}\left[Le^{-(T-t)(a+q)L}\left(I+cL\int_{t}^{T}\eta_{s}e^{-(T-s)(a+q)L}\,\mathrm{d}s\right)^{-1}\right]\\ \left[\det\left(I+cL\int_{t}^{T}\eta_{s}e^{-(T-s)(a+q)L}\,\mathrm{d}s\right)\right]^{\frac{1}{N}}.$ (4.48)
Observe that equation (4.48) involves only the model parameters and the graph Laplacian $L$ without relying on $P_{t}$, which aligns with our goal of characterizing $\eta_{t}$. Nevertheless, equation (4.48) remains complex and does not easily allow for proving existence and uniqueness. To address this, we apply matrix calculus techniques. The lemma below plays an important role in formulating the governing equation for $\eta_{t}$.
Lemma 4.9 ([PP08, Equation (41)]).
$\frac{\mathrm{d}\det A(t)}{\mathrm{d}t}=\emph{\text{Tr}}\left(\emph{\text{adj}}[A(t)]\frac{\mathrm{d}A(t)}{\mathrm{d}t}\right),$ (4.49)
where $\emph{\text{adj}}B=B^{-1}\det B$ is the adjugate matrix, well-defined for any nonsingular matrix $B$.
$\frac{\mathrm{d}\det A(X)}{\mathrm{d}X}=\frac{\emph{\text{Tr}}\left(\emph{\text{adj}}[A(X)]\mathrm{d}A(X)\right)}{\mathrm{d}X}.$ (4.50)
The next corollary follows easily from Lemma 4.50 and the chain rule of matrix calculus.
Corollary 4.10.
$\frac{\mathrm{d}[\det(I+cLR(t))]^{\frac{1}{N}}}{\mathrm{d}t}=\frac{1}{N}\emph{\text{Tr}}\left([I+cLR(t)]^{-1}cLR^{\prime}(t)\right)[\det(I+cLR(t))]^{\frac{1}{N}}.$ (4.51)
$\frac{\mathrm{d}[\det(I+cXL)]^{\frac{1}{N}}}{\mathrm{d}X}=\frac{1}{N}[\det(I+cXL)]^{\frac{1}{N}}(I+cXL)^{-1}cL.$ (4.52)
$R(t):=\int_{T-t}^{T}\eta_{s}e^{-(T-s)(a+q)L}\,\mathrm{d}s,$ (4.53)
then, based on equation (4.48) and Corollary 4.10 and the definition of $Q$, $R(t)$ solves the following matrix-valued ODE:
$R^{\prime}(T-t)=\eta_{t}e^{-(T-t)(a+q)L}$ (4.55)
provides the corresponding value of $\eta_{t}$.
Remark 4.11 (Assumptions for heuristic derivations).
The following assumptions are made in the heuristic construction:
  1. (A.1)
  2. (A.2)
  3. (A.3)
    $\Lambda^{i}_{t}$ is invertible, $\forall t\in[0,T]$, $\forall i\in[N]$.
  4. (A.4)
4.2.4 The construction of the Markovian NE
Once $\eta_{t}$ has been constructed based on equations (4.3)–(4.4), the Markovian NE can be determined as follows. From equations (4.37) and (4.55), $P_{t}$ can be represented in terms of $R$ as follows:
$P_{t}=(a+q)L+R^{\prime}(T-t)\;cL[I+R(T-t)\;cL]^{-1}.$ (4.56)
which coincides with equation (4.6). By combining equations (4.29)–(4.30) and (4.39), we obtain
$F^{i}_{t}=\frac{c}{1-c\int_{t}^{T}\left(\frac{\text{Tr}(\tilde{P}(s)L)}{N}\right)^{2}\,\mathrm{d}s}\tilde{P}_{t}Le_{i}e_{i}^{\operatorname{T}}L\tilde{P}_{t}.$ (4.57)
The denominator of $F^{i}_{t}$ can be simplified using equations (4.41)–(4.42):
$1-c\int_{t}^{T}\left(\frac{\text{Tr}(\tilde{P}_{s}L)}{N}\right)^{2}\,\mathrm{d}s=\frac{1}{cN\eta^{2}_{t}}[\text{Tr}(P_{t})-(a+q)\text{Tr}(L)].$ (4.58)
Plugging equations (4.33) and (4.58) into equation (4.57) leads to the result (4.5).
@@ 4.2.4 The construction of the Markovian NE
There exists $t_{0}\in\mathbb{R}_{+}$ and $x_{0}\in(0,\frac{1}{4c})$ such that the solution $R(t)$ to equation (4.3) exists and is unique in $\mathscr{X}$, when $t\in[0,t_{0}]$ and $0\leq R(t)\leq x_{0}I$.

### v2

@@ 4 Semi-Explicit Equilibrium under Vertex-Transitive Graph
$\text{Aut}(G)\hskip-2.0pt:=\hskip-2.0pt\{\text{bijection }\phi:V\to V|(u,v)\in E\text{ if and only if }(\phi(u),\phi(v))\in E,\ \forall u,v\in V\}.$ (4.1)
In the definition of $\text{Aut}(G)$, a graph automorphism $\phi$ should be understood as a bijective relabeling of graph vertices that preserves edge connections. Algebraically, $\text{Aut}(G)$ has a group structure under the operation of function composition, which induces a natural group action on the vertex set $V$ via $\text{Aut}(G)\times V\ni(\phi,v)\mapsto\phi(v)\in V$. From an algebraic perspective, a graph is called vertex-transitive if such group action on the vertex set is transitive, i.e., all graph vertices are indistinguishable and structurally equivalent.
The strong symmetry of vertex-transitive graphs facilitates theoretical analysis and makes them appealing for practical applications. For example, in terms of graph homomorphisms, the graph cores of vertex-transitive graphs have special structures [GR01]. In terms of spectral properties, tighter spectral bounds can be established for vertex-transitive graphs [Chu97]. On the application side, vertex-transitive graphs are crucial in communication network designs [TA92].
@@
where the inequality between symmetric matrices is defined in terms of the semi-positive-definite sense. That is, $A\geq B$ if and only if ABA-B is semi-positive-definite, for $A,B\in\mathbb{S}^{N\times N}$.
@@ 4.1 Main results
$F^{i}_{t}=\frac{1}{\frac{\text{Tr}(P_{t})-(a+q)\text{Tr}(L)}{N}}[P_{t}-(a+q)L]e_{i}e_{i}^{\operatorname{T}}[P_{t}-(a+q)L],$ (4.5)
@@
Equations (4.3)–(4.6) provide an explicit construction of a global solution to the Riccati system (2.14). By [Luk71, Theorem 2.3 and Section 2.4], the solution to the coupled system (2.14) is unique whenever it exists. Combining both results establishes the global well-posedness of system (2.14) when $G$ is vertex-transitive and $\varepsilon=q^{2}$. This nontrivial result is presented in Corollary 4.3.
Corollary 4.3.
Let $G$ be a simple connected vertex-transitive graph with undirected edges. Assume $q^{2}=\varepsilon$ in problem (2.2)–(2.4). Then for any $T>0$, the coupled Riccati system (2.14) admits a unique solution on $[0,T]$, given by equations (4.3)–(4.6).
Compared to Corollary 3.11, which only provides an existence interval for the solution, Corollary 4.3 provides a much stronger result, thanks to the adequate graph symmetry and the uncoupled structure of model parameters (see Remark A.3).
We note that the existence of a global solution to the coupled Riccati system (2.14) has not yet been fully characterized in the existing literature. While some studies have approached the solvability of coupled Riccati equations from a geometric perspective [PO84], global existence for any $T>0$ and $N\geq 2$ is only guaranteed when all coefficients are diagonal matrices (see [AKFIJ12, part (iii) of Remark 6.6.3]), which means that system (2.14) requires a trivial diagonal graph Laplacian $L$. By exploiting symmetry and decoupling the equations within the Riccati system (2.14), Corollary 4.3 greatly improves existing global well-posedness results.
@@
where $P_{t}$ takes the form
$P_{t}=\sum_{k=1}^{N}\left[(a+q)L+F^{k}_{t}\right]e_{k}e_{k}^{\operatorname{T}}=(a+q)L+\sum_{k=1}^{N}F^{k}_{t}e_{k}e_{k}^{\operatorname{T}},$
and has the representation (4.6) given by Theorem 4.2. Given the initial condition $\hat{X}_{0}=X_{0}$, this Ornstein-Uhlenbeck dynamics has a closed-form solution:
@@
where we have used the fact that $P_{s}$ and $P_{t}$ commute for any $t,s\in[0,T]$. Therefore, $\hat{X}_{t}$ is Gaussian distributed with mean $e^{-\int_{0}^{t}P_{s}\,\mathrm{d}s}X_{0}$ and variance $\sigma^{2}\int_{0}^{t}e^{-2\int_{s}^{t}P_{u}\,\mathrm{d}u}\,\mathrm{d}s$. Using equations (A.16) and (A.34) in Appendix A below, which can be verified to be correct, the mean and variance of $\hat{X}_{t}$ can be represented in terms of $R$, the solution to equation (4.3).
We note the following, (i) Theorem 4.2 applies specifically to the case where $q^{2}=\varepsilon$. As will discussed in Remark A.3, this condition results in an uncoupled nature of the underlying linear system, greatly simplifying the calculations. (ii) A numerical verification of Theorem 4.2 will be presented in Section 5.1. (iii) Theorem 4.2 can be generalized to provide a semi-explicit Markovian NE for games where the mean reversion level is the average of states within a more broadly defined neighborhood of a player, such as the $l$-neighborhood (see Section 4.1.1). This extends the concept of a graph neighborhood as introduced in Definition 2.1. The generalized result is briefly outlined below.
4.1.1 Generalizations to $l$-neighborhood interactions
@@
To set the mean reversion level as the average within the $l$-neighborhood, we consider the following generalized state dynamics on a vertex-transitive graph $G$:
@@
$f^{i}(t,x,\alpha)$ $=\frac{1}{2}(\alpha)^{2}-q\alpha\Big{(}\sum_{j:v_{j}\in N_{G}^{l}(v_{i})}\frac{n(v_{j},v_{i};l)}{|N_{G}^{l}(v_{i})|}x^{j}-x^{i}\Big{)}$ (4.14)
$\qquad+\frac{\varepsilon}{2}\Big{(}\sum_{j:v_{j}\in N_{G}^{l}(v_{i})}\frac{n(v_{j},v_{i};l)}{|N_{G}^{l}(v_{i})|}x^{j}-x^{i}\Big{)}^{2},$ (4.15)
$g^{i}(x)$ $=\frac{c}{2}\Big{(}\sum_{j:v_{j}\in N_{G}^{l}(v_{i})}\frac{n(v_{j},v_{i};l)}{|N_{G}^{l}(v_{i})|}x^{j}-x^{i}\Big{)}^{2}.$ (4.16)
@@
where each vertex $v_{j}\in\bigcup_{k=1}^{l}N^{k}_{G}(v_{i})$ is assigned a weight proportional to a discounted sum of the number of paths of length $k\in\left\{1,\ldots,l\right\}$ between $v_{i}$ and $v_{j}$. Each path of length $k$ is discounted by a factor of $w^{k}$, where $w$ denotes the discount rate.
The expression in equation (4.19) can be written in matrix form $e_{i}^{\operatorname{T}}S(l,L)X_{t}$, where
$S(l,L)=I-\frac{1-w\delta}{1-(w\delta)^{l}}(I-L)[I-(w\delta)^{l}(I-L)^{l}][I-w\delta(I-L)]^{-1}.$ (4.20)
@@
4.1.2 Proof roadmap
This subsection briefly outlines the key steps leading to formulation (4.5), with further details provided in Appendix A.
@@ 4.1.2 Proof roadmap
The construction of (4.5), together with (4.3)–(4.4) and (4.6), is inspired by a previous work [LS22] and employs two key techniques: a fixed point scheme that helps represent the solution to the Riccati system, and the regular representation of the graph automorphism group that leverages symmetry.
Step 1: We first define the quantity
$P_{t}:=\sum_{k=1}^{N}\left[(a+q)L+F^{k}_{t}\right]e_{k}e_{k}^{\operatorname{T}}=(a+q)L+\sum_{k=1}^{N}F^{k}_{t}e_{k}e_{k}^{\operatorname{T}},$
which appears in system (2.14). Solving for $F^{i}$ in terms of $P$ yields (A.10).
Step 2: Next, we substitute equation (A.10) into the above definition of $P$, introduce a new quantity $\eta$ (see equation (A.12)), and express $P$ in terms of $\eta$, leading to equation (A.18).
Step 3: Using the definition of $\eta$, we derive the governing equation:
$R^{\prime}(T-t)=\eta_{t}e^{-(T-t)(a+q)L},$
where $R$ solves equation (4.3).
Step 4: Once $R$ is obtained from the matrix-valued ODE (4.3), and $\eta$ is constructed, we can represent $P_{t}$ in terms of $R$, rather than in terms of $\eta$, leading to (4.6). Finally, combining equations (4.6) and (A.10) gives (4.5).
The full details can be found in Appendices A.1A.4. The following two subsections take (4.5) as given and focus on establishing the well-posedness of equation 4.3, which ensures the existence of $\eta$, and on verifying that equation (4.5) indeed solves the Riccati system (2.14).
@@
There exists $t_{0}\in\mathbb{R}_{+}$ and $x_{0}\in\mathbb{R}_{+}$ such that the solution $R(t)$ to equation (4.3) exists and is unique in $\mathscr{X}$, when $t\in[0,t_{0}]$ and $0\leq R(t)\leq x_{0}I$.